Activity One: Discrete Probability Distributions in Action (Example #1)

A company examining its website operations notes the number of overloads being experienced follows a Poisson distribution with an average of three overloads to the online- booking system during a 100 day period. Express your answers in a percentage to two decimal places, what is the probability...

Q1.1) Of no failures on a given day?
Q1.2) One or more overloads on a given day?

Q1.3) At least two failures in a three day period?
Answer #1

Here parameter of poisson distribution \lambda = 3

if x is the number of failures for the overloads.

Question 1.1

Pr(x = 0) = e-3 = 0.0498

Question 1.2

Pr(x >= 1) = 1 - e-3 = 0.9502

Question 1.3

Pr(x >= 2) = 1 - e-3 - 3 e-3 = 0.8009